OpenStudy (albert0898):

In triangle ABC, a = 2x, b = 3x + 2, c = radical(12), and angle C = 60. Find x.

OpenStudy (albert0898):



OpenStudy (albert0898):

I know that I have to use the law of cosines, but I keep getting a different answer from the book.......

OpenStudy (serenity74):

Can I see your work?

OpenStudy (albert0898):

I just plugged in the sides and angle in the following formula \[a^2 = b^2 + c^2 - 2bc \cos A\]

OpenStudy (serenity74):

Right formula but what did you plug into the letters?

OpenStudy (albert0898):

\[(\sqrt{12})^2 = (2x)^2 + (3x+2)^2 - 2(2x)(3x+2) * \cos 60\]

OpenStudy (serenity74):

Wrong formula actually. Cos 60 is c's NOT a's

OpenStudy (albert0898):

That's the general formula

OpenStudy (albert0898):

You can flip the formula around

OpenStudy (serenity74):

IK but you have to flip it around first and then solve for C^2

OpenStudy (serenity74):

and then you use sine formula to make life easier

OpenStudy (albert0898):

By flip around, I mean you can switch the letters. Instead of a, you can b or c. It's just a general formula that can be changed around to best suit the problem.

OpenStudy (serenity74):

IKIKIKIKIKIK

OpenStudy (serenity74):

Did you do that doe? And you still didn't get the right answer? Did you check the textbook's work? You know what let me work it out on a piece of paper and I will ask you what you got okay?

OpenStudy (serenity74):

if we got the same answer then the book is wrong okay?

OpenStudy (albert0898):

I did do that. The textbook did it the same way I did it, but somehow got the wrong answer. I can type up the textbook's explanation if you want?

OpenStudy (serenity74):

Yes please? I am so sorry for bombarding you with tonz of questions dude.

OpenStudy (serenity74):

@Albert0898 I will be back in 30 minutes or so cause I have to drive back home so I will give you your answer as soon as i get home so sorry!!!

OpenStudy (albert0898):

It's ok haha Here's the textbook's explanation. --------- Law of Cosines: \[12 = (2x)^2 + (3x+2)^2 - 2(2x)(3x+2) \cos60\] \[12 = 4x^2 + 9x^2 + 12x + 4 - (12x^2 + 8x) \frac{ 1 }{ 2 }\] \[7x^2 + 8x - 8 = 0\] Use program QUADFORM to get \[x = pm0.64\] Since a side of a triangle must be positive, x can only equal 0.64.

imqwerty (imqwerty):

yeah thats correct and this is the solution that u'll be getting by solving your equation u might be doing some calculation mistake check again :)

OpenStudy (albert0898):

Okay @imqwerty In this book, it does not divide both sides by cos 60. That's where my mistake was. In my ACT Textbooks, it would say that I would have to do that. I'm very confused. I see my mistake now, but am still very confused as to why two different books have two different methods that yield two very different answers.

imqwerty (imqwerty):

:) we know that cos 60 is \(\large \frac{1}{2}\) so just simply put it there in the formula and rest of simplification is normal.

OpenStudy (albert0898):

But why is it that when I divide 1/2 from both sides, it doesn't work out to get the right answer?

imqwerty (imqwerty):

it must work alright can you show your work?

OpenStudy (albert0898):

Sure thing! \[12 = (4x)^2 + (9x)^2 + 12x + 4 - (12x^2 + 8x)(\frac{ 1 }{ 2 })\] \[24 = 13x^2 + 12x + 4 - (12x^2 + 8x)\] \[24 = x^2 + 4x + 4\] \[x^2 + 4x - 20 = 0\]

imqwerty (imqwerty):

1st step has a lil mistake-> \(12 = (4x^{\color{red}2})^{\color{red}1} + (9x^{\color{red}2})^{\color{red}1} + 12x + 4 - (12x^2 + 8x)(\frac{ 1 }{ 2 })\) and then when u divide by \(\large \frac{1}{2}\) it goes like this- \(\large\frac{12}{\frac{1}{2}} = \frac{13x^2+12x + 4 - (12x^2 + 8x)(\frac{ 1 }{ 2 })}{\frac{1}{2}}\) \(24 = \large{\frac{13x^2 + 12x + 4}{\color{red}{\frac{1}{2}}}} - (12x^2 + 8x)\) did you see the fault? :)

OpenStudy (albert0898):

Yes, I did think of that as my fault. However, my ACT books (which I don't use anymore) wouldn't state that as a method. It might have been a typo though

imqwerty (imqwerty):

its good to directly substitute the value rather than dividing because dividing will make the process lengthy in most cases

OpenStudy (albert0898):

Okay, well I learned something new today! Thank you @imqwerty ! :)

imqwerty (imqwerty):

np :)

OpenStudy (serenity74):

welp nvm den sorry!